Katherine 12m

Disk VSN: HOB+0042/2000 | Data volume at beginning: 8210.235 GB

  • 18:30 Experiment started OK. (Ross)
  • 07:33 UT dbbc drifting. Reconfiguring the dbbc. Affected scans: 271-0737, 271-0738b, 271-0742 (Simin).

Yarragadee 12m

Disk VSN: NAIC-004/2000 | Data volume at beginning: 0.000 GB

  • 18:30 Experiment started OK. (Ross)
  • 15:10UT Noticed the System Monitor for Yg says antenna is stowed, however Az El are changing and no indication of stow on timeyg. Maybe something for Jamie to check? (Arwin)